OpenStudy (anonymous):

A rectangle has a length of x + 2, a width of 7, and a perimeter of 32. The value of x is _______

OpenStudy (anonymous):

medal me first for it !#

OpenStudy (gorv):

perimeter =sum of length of all sides

OpenStudy (anonymous):

okay if you give the right answer..

OpenStudy (anonymous):

x=7

OpenStudy (gorv):

x+2+7+x+2+7=32 2x+18=32 2x=32-18=14 x=14/2=7

OpenStudy (anonymous):

The value of x isnt 14

OpenStudy (anonymous):

The value of X is 7

OpenStudy (anonymous):

soo gorv was wrong?